OpenStudy (anonymous):

Looking for a tutor... A software designer is mapping the streets for a new racing game. All of the streets are depicted as either perpendicular or parallel lines. The equation of the lane passing through A and B is -7x + 3y = -21.5. What is the equation of the central street PQ?

OpenStudy (anonymous):

This is the diagram.

OpenStudy (anonymous):

The choices are A) -3x + 4y = 3 B) -1.5x − 3.5y = -31.5 C) 2x + y = 20 D) -2.25x + y = -9.75

OpenStudy (anonymous):

Never mind. To anyone who gets this in the future, it's going to be B. I took a wild guess and got it right. :/